OpenStudy (anonymous):

Solve: -1/2 x + 3 = −x + 7

OpenStudy (anonymous):



OpenStudy (phi):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Something Wrong?

OpenStudy (phi):

what does -1/2 mean?

OpenStudy (anonymous):

I don't know if there using -1/2 as a fraction or division I think fraction but my schools stupid so I drew how they put it.

OpenStudy (anonymous):

You never know with my school

OpenStudy (phi):

Is the problem \[ - \frac{1}{2} x +3 = -x+7\]

OpenStudy (anonymous):

Yes! Lol sorry i'm not good at math at all....

OpenStudy (phi):

As a first step, I would add +x to both sides like this \[x - \frac{1}{2} x +3 = x -x+7\] and simplify it

OpenStudy (anonymous):

Okay? Still a little confused sorry.

OpenStudy (phi):

can you simplify the right side: x-x+7

OpenStudy (anonymous):

Umm no

OpenStudy (phi):

x-x +7 means you start with 1 x, you take away 1 x , and add 7 what do you get?

OpenStudy (anonymous):

8x? or does it stay 7?

OpenStudy (phi):

think of x a "real thing" you have x's and you have numbers. x -x is 0 (x take away one x is 0) 0+7 is 7 so now you have \[ x - \frac{1}{2} x +3 = 7\] on the left you have 1 x and take way 1/2 of an x. How much of x do you now have?

OpenStudy (anonymous):

1/2?

OpenStudy (phi):

1/2 x (don't forget the x) the equation is now ,after we replace x - 1/2 x with 1/2 x \[ \frac{1}{2}x +3 = 7\] I would now add -3 to both sides. write -3 on both sides. can yo do that?

OpenStudy (anonymous):

Sorry my computer froze

OpenStudy (phi):

I would now add -3 to both sides. write -3 on both sides. can yo do that?

OpenStudy (anonymous):

Sorry my computer keeps messing up -3 + 1/2x +3 = 7 + 3-

OpenStudy (phi):

ok, except people generally don't write -3 as 3- -3 + 1/2x +3 = 7 + -3 on the left side we have 3 "terms" we can re-order them (Example: 1+2+3 is the same as 2+3 + 1) 1/2 x + 3 + -3 = 7 + -3 what is 3 take away 3? what is 7 take away 3 ? can you re-write the equation after simplifying?

OpenStudy (anonymous):

0 and 4

OpenStudy (phi):

yes, Can you write down what the equation looks like?

OpenStudy (anonymous):

1/2x + 4

OpenStudy (phi):

try again. equations have = signs in them

OpenStudy (phi):

1/2 x + 3 + -3 = 7 + -3 simplifies to 1/2 x = 4

OpenStudy (anonymous):

1/2x =4

OpenStudy (phi):

last step, multiply both sides by 2 (if we double both sides, the equation is still true) 2* 1/2 * x = 4*2 (* means multiply here) now simplify

OpenStudy (anonymous):

I got 8 for 4*2 but how do I multiply the fraction? Yeah I know I should know how to do it...

OpenStudy (phi):

when you have time look at http://www.khanacademy.org/math/arithmetic/fractions and look for multiplying (or anything else you need to brush up on.) to multiply fractions you multiply top times top and bottom times bottom if you have a whole number, you assume its bottom is 1 so 2*1/2 is \[ \frac{2}{1} \cdot \frac{1}{2} \] can you finish?

OpenStudy (anonymous):

2/2

OpenStudy (phi):

yes, and 2 divided by 2 is ?

OpenStudy (anonymous):

1

OpenStudy (phi):

that means we simplify 2* 1/2 * x = 4*2 to 1 x =8 or (because you don't bother writing 1 in front of the x) x= 8 the answer.

OpenStudy (anonymous):

Thank you

OpenStudy (phi):

Here are videos on how to solve equations. http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities

OpenStudy (anonymous):

Okay Thank you for helping me. I'll look at the site.